# Bounds on the realizations of zero-nonzero patterns and sign conditions of polynomials restricted to varieties and applications

### v2

@@ 4.8. Proofs of Theorems 12 and 13
For $1\leq t\leq t_{0}$, let $V_{t}=A^{t}$, and let $V=\oplus_{1\leq t\leq t_{0}}V_{t}$ denote the direct sum and $i_{t}:V_{t}\rightarrow V$ the canonical injection. Let $\Delta^{\prime}=\cup_{1\leq t\leq t_{0}}i_{t}(\Sigma^{t})$, noting that $\Delta^{t}\subset V_{t}$ is an algebraic subset of $V_{t}$.
Notice that $\Delta^{\prime}$ is an algebraic subset of dimension bounded by $pt_{0}$ and degree bounded by $D+D^{2}+\cdots+D^{t_{0}}\leq D^{t_{0}+1}$. Finally, let $\phi:V\rightarrow A$ denote the map induced (by extending linearly) the maps $\phi_{t}:V_{t}\rightarrow A$, sending $(v_{1},\ldots,v_{t})$ to $v_{1}+\cdots+v_{t}$ (i.e. $\phi=\sum_{1\leq t\leq t_{0}}\phi_{t}\circ\pi_{t}$, where $\pi_{t}:V\rightarrow V_{t}$ is the canonically defined projection).
@@
For $1\leq t\leq t_{0}$, let $V^{*}=\oplus_{1\leq t\leq t_{0}}V^{t}$ denote the direct sum and $i_{t}:V^{t}\rightarrow V$ the canonical injection. Let $\Delta^{\prime}=\cup_{1\leq t\leq t_{0}}i_{t}(\Delta^{t})$, noting that $\Delta^{t}\subset UV^{t}L$ is a real algebraic subset of $V^{t}$.
Notice that $\Delta^{\prime}$ is a union of non-singular complete intersection subvarieties of dimension bounded by $pt_{0}$ and degree bounded by $D+D^{2}+\cdots+D^{t_{0}}\leq D^{t_{0}+1}$. Finally, as before let $\phi:V^{*}\rightarrow V$ denote the map induced (by extending linearly) the maps $\phi_{t}:V^{t}\rightarrow V$, sending $(v_{1},\ldots,v_{t})$ to $v_{1}+\cdots+v_{t}$ (i.e. $\phi=\sum_{1\leq t\leq t_{0}}\phi_{t}\circ\pi_{t}$, where $\pi_{t}:V^{*}\rightarrow V^{t}$ is the canonically defined projection).
@@ 4.9. Proofs of Theorems 14 and 15
For $1\leq t\leq t_{0}$, let $V_{t}=A^{t}$, and let $V=\oplus_{1\leq t\leq t_{0}}V_{t}$ denote the direct sum and $i_{t}:V_{t}\rightarrow V$ the canonical injection. Let $\Delta^{\prime}=\cup_{1\leq t\leq t_{0}}i_{t}(\Sigma^{t})$, noting that $\Delta^{t}\subset V_{t}$ is an algebraic subset of $V_{t}$.
Notice that $\Delta^{\prime}$ is an algebraic subset of dimension bounded by $pt_{0}$ and degree bounded by $D+D^{2}+\cdots+D^{t_{0}}\leq D^{t_{0}+1}$. Finally, let $\phi:V\rightarrow A$ denote the map induced (by extending linearly) the maps $\phi_{t}:V_{t}\rightarrow A$, sending $(f_{1},\ldots,f_{t})$ to $f_{1}\cdots f_{t}$ (i.e. $\phi=\sum_{1\leq t\leq t_{0}}\phi_{t}\circ\pi_{t}$, where $\pi_{t}:V\rightarrow V_{t}$ is the canonically defined projection).
@@
We follow the proof of Theorem 14. If $\mathrm{mmrank}_{A,\Delta}(W)=\infty$, then there is nothing to prove. Otherwise, let $\mathrm{mmrank}_{A,\Delta}(W)=t_{0}$.
For $1\leq t\leq t_{0}$, let $V_{t}=A^{t}$, and let $V=\oplus_{1\leq t\leq t_{0}}V_{t}$ denote the direct sum and $i_{t}:V_{t}\rightarrow V$ the canonical injection. Let $\Delta^{\prime}=\cup_{1\leq t\leq t_{0}}i_{t}(\Delta^{t})$, noting that $\Delta^{t}\subset V_{t}$ is a real algebraic subset of $V_{t}$.
Notice that $\Delta^{\prime}$ is a union of non-singular complete intersection subvarieties of dimension bounded by $pt_{0}$ and degree bounded by $D+D^{2}+\cdots+D^{t_{0}}\leq D^{t_{0}+1}$. Finally, as before let $\phi:V\rightarrow A$ denote the map induced (by extending linearly) the maps $\phi_{t}:V_{t}\rightarrow A$, sending $(f_{1},\ldots,f_{t})$ to $f_{1}\cdots f_{t}$ (i.e. $\phi=\sum_{1\leq t\leq t_{0}}\phi_{t}\circ\pi_{t}$, where $\pi_{t}:V\rightarrow V_{t}$ is the canonically defined projection).
@@ 4.10. Proofs of Theorems 16 and 17
Let $\Delta=\bigcup_{\mathbf{i}=(i_{1},\ldots,i_{q}),1\leq i_{1}<\cdots<i_{q}\leq n+1}A_{\mathbf{i}}$.
@@
It is easy to set that $W^{+}$ is a $\mathcal{P}$-semi-algebraic subset of $A$, for some $\mathcal{P}$ with $\mathrm{card}(\mathcal{P}\leq O(2^{2(n+t)})$ and $\deg(P)\leq 1$.
@@ 5. Open problems and future work
  3. 3.
    We believe that the ambient dimension free upper bounds proved in this paper should have other applications (especially in computational complexity theory) than those explored in this paper. We would like to explore such applications.

### v3

@@ 4.7. Proofs of Theorems 10 and 11
If $\mathrm{rank}_{V,\Delta}(W)=\infty$, then there is nothing to prove. Otherwise, let $\mathrm{rank}_{V,\Delta}(W)=t_{0}$.
For $1\leq t\leq t_{0}$, let
$V^{t}=V\underbrace{\times\cdots\times}_{\text{$t$ times}}V,$
and let
$V^{\prime}=\bigoplus_{1\leq t\leq t_{0}}V^{t}$
denote the direct sum and
$i_{t}:V^{t}\rightarrow V^{\prime}$
the canonical injection.
Note that $\Delta^{t}=\Delta\underbrace{\times\cdots\times}_{\text{$t$ times}}\Delta$ is an algebraic subset of $V^{t}$ with $\dim(\Delta^{t})\leq pt$ and $\deg\Delta^{t}\leq D^{t}$.
Let
$\Delta^{\prime}=\bigcup_{1\leq t\leq t_{0}}i_{t}(\Delta^{t}).$
Notice that $\Delta^{\prime}$ is an algebraic subset of $V^{\prime}$ with $\dim\Delta^{\prime}\leq pt_{0}$ and $\deg\Delta^{\prime}\leq D+D^{2}+\cdots+D^{t_{0}}\leq D^{t_{0}+1}$. Finally, let $\phi:V^{\prime}\rightarrow V$ denote the map induced (by extending linearly) the maps $\phi_{t}:V^{t}\rightarrow V$, sending $(v_{1},\ldots,v_{t})$ to $v_{1}+\cdots+v_{t}$ (i.e. $\phi=\sum_{1\leq t\leq t_{0}}\phi_{t}\circ\pi_{t}$, where $\pi_{t}:V^{\prime}\rightarrow V_{t}$ is the canonically defined projection).
@@
For $1\leq t\leq t_{0}$, let
$V^{\prime}=\bigoplus_{1\leq t\leq t_{0}}V^{t}$
denote the direct sum and $i_{t}:V^{t}\rightarrow V$ the canonical injection. Let
$\Delta^{\prime}=\bigcup_{1\leq t\leq t_{0}}i_{t}(\Delta^{t}),$
noting that $\Delta^{t}\subset V^{t}$ is a real algebraic subset of $V^{t}$.
Notice that $\Delta^{\prime}$ is a union of non-singular complete intersection subvarieties of dimension bounded by $pt_{0}$ and degree bounded by $D+D^{2}+\cdots+D^{t_{0}}\leq D^{t_{0}+1}$. Finally, as before let $\phi:V^{\prime}\rightarrow V$ denote the map induced (by extending linearly) the maps $\phi_{t}:V^{t}\rightarrow V$, sending $(v_{1},\ldots,v_{t})$ to $v_{1}+\cdots+v_{t}$ (i.e. $\phi=\sum_{1\leq t\leq t_{0}}\phi_{t}\circ\pi_{t}$, where $\pi_{t}:V^{\prime}\rightarrow V^{t}$ is the canonically defined projection).
@@ 4.8. Proofs of Theorems 12 and 13
For clarity we denote by $V$ the underlying vector space of $A$.
For $1\leq t\leq t_{0}$, let
$V^{t}=V\underbrace{\times\cdots\times}_{\text{$t$ times}}V,$
and let
$V^{\prime}=\bigoplus_{1\leq t\leq t_{0}}V^{t}$
denote the direct sum and $i_{t}:V_{t}\rightarrow V^{\prime}$ the canonical injection. Let
$\Delta^{\prime}=\cup_{1\leq t\leq t_{0}}i_{t}(\Delta^{t}),$
noting that $\Delta^{t}\subset V_{t}$ is an algebraic subset of $V_{t}$.
Notice that $\Delta^{\prime}$ is an algebraic subset of dimension bounded by $pt_{0}$ and degree bounded by $D+D^{2}+\cdots+D^{t_{0}}\leq D^{t_{0}+1}$. Finally, let $\phi:V^{\prime}\rightarrow V$ denote the map induced (by extending linearly) the maps $\phi_{t}:V_{t}\rightarrow V$, sending $(f_{1},\ldots,f_{t})$ to $f_{1}\cdots f_{t}$ (i.e. $\phi=\sum_{1\leq t\leq t_{0}}\phi_{t}\circ\pi_{t}$, where $\pi_{t}:V^{\prime}\rightarrow V_{t}$ is the canonically defined projection).
@@
We follow the proof of Theorem 12 and follow the notation inroduced there. If $\mathrm{mmrank}_{A,\Delta}(W)=\infty$, then there is nothing to prove. Otherwise, let $\mathrm{mmrank}_{A,\Delta}(W)=t_{0}$.
Notice that in this case $\Delta^{\prime}$ (see definition in proof of Theorem 12) is a union of non-singular complete intersection subvarieties of dimension bounded by $pt_{0}$ and degree bounded by $D+D^{2}+\cdots+D^{t_{0}}\leq D^{t_{0}+1}$. Finally, as before let $\phi:V\rightarrow A$ denote the map induced (by extending linearly) the maps $\phi_{t}:V_{t}\rightarrow A$, sending $(f_{1},\ldots,f_{t})$ to $f_{1}\cdots f_{t}$ (i.e. $\phi=\sum_{1\leq t\leq t_{0}}\phi_{t}\circ\pi_{t}$, where $\pi_{t}:V\rightarrow V_{t}$ is the canonically defined projection).
@@
$\sum_{\sigma\in\{0,1,-1\}^{\mathcal{P}}}b_{0}(\phi^{-1}({\mathcal{R}}(\sigma))\cap\Delta^{\prime})\leq D^{2(t_{0}+1)}(t_{0}s\delta)^{O(pt_{0})}.$
@@ 4.9. Proofs of Theorems 14 and 15
Let
$\Delta=\bigcup_{\mathbf{i}=(i_{1},\ldots,i_{q}),1\leq i_{1}<\cdots<i_{q}\leq n+1}A_{\mathbf{i}}.$
@@
It is easy to set that $W^{+}$ is a $\mathcal{P}$-semi-algebraic subset of $A$, for some $\mathcal{P}$ with $\mathrm{card}(\mathcal{P})=O(2^{2(n+t)})$ and $\deg(P)\leq 1$.